2-(4y-x/2)=5x+5


Need the slope?

Answer by ptth176(50)

2-(4y-x/2)=5x+5
=>2-4y+x/2=5x+5
=>-3-(9/2)x=4y
=>y=(-9/8)x-3/4
slope:m=-9/8